# SparseVLM: Visual Token Sparsification for Efficient Vision-Language Model Inference

### v2

@@ 4.2 Video Understanding Tasks
Datasets. We test our method on four common video question answering benchmarks, TGIF-QA (Jang et al., 2017), MSVD-QA (Xu et al., 2017), MSRVTT-QA (Xu et al., 2017) and ActivityNet-QA (Yu et al., 2019), where video-question pairs are massively disproportional in length. We adopt the evaluation framework proposed by Video-ChatGPT (Maaz et al., 2023) that utilizes both accuracy and ChatGPT score as key performance metrics with details in the Appendix A.4.
Implementation Details. We directly apply our SparseVLM for Video-LLaVA (Lin et al., 2023), which is a commonly used VLM framework for video question answering. Video-LLaVA is composed of several key components, including language bind encoder $f^{v}_{M}$(Zhu et al., 2023a) for extracting features from raw visual inputs (e.g., images or videos), a language decoder model $f_{L}$ such as Vicuna (Touvron et al., 2023), a visual projection layer $f_{P}$, and a word embedding layer $f_{T}$. We adopt the same inference setup as the original Video-LLaVA code base222github.com/PKU-YuanGroup/Video-LLaVA., as it is publicly available.
Main Results. In Table 2, we set the Video-LLaVA with $2048$ video tokens as our upper bound for an overall average accuracy of $100.0\%$ and a score of $+0.00$. To make a fair comparison, we both preserve $135$ vision tokens ($93.4\%$ pruning ratio) for FastV (Chen et al., 2024b) and SparseVLM. It is clear that our approach consistently outperforms FastV across all benchmarks, both in accuracy (Acc.) and GPT evaluation score. SparseVideoLLaVA achieves a total average accuracy of $86.5\%$, a significant 34.4% higher than $52.1\%$ of FastV. From the GPT score perspective, SparseVLM only loses $0.17$ points compared to $1.02$ points of FastV. These improvements suggest that when handling video modality containing temporal features, SparseVLM continues to deliver strong performance, generating accurate responses to diverse questions while utilizing significantly fewer tokens. This achieves an effective trade-off between inference efficiency and model performance.
5 Analysis
5.1 Effects of Relevant Text Token Selection
Figure 5: The ablation study of text raters.
We propose a selection mechanism to localize visually irrelevant text tokens to limit their negative effects in rating the significance of vision tokens. Here we conduct experiments to analyze the effects of the mechanism in Figure 5. Under the same number of vision tokens ($142$), we have 3 settings (using all tokens, only text tokens, and only text raters we select) with LLaVA (Liu et al., 2023a) to judge vision token candidates. In TextVQA (Singh et al., 2019), our mechanism improves the vanilla text-aware method (only text tokens) by $0.79\%$, which validates that our extra selection is effective. Besides, by building upon the text-aware manner, we further outperform the baseline (all tokens) by $\textbf{4.3}\%$ on POPE (Li et al., 2023b). The huge margin means POPE sparsification is quite sensitive to question prompts, and text guidance is necessary. In summary, text rater selection is general and improves the performance across various scenarios.
5.2 Effects of Pruned Tokens Recycling
Table 3: Ablation study on token reconstruction (TR). Experiments are conducted on TextVQA and POPE using LLaVA with various sparsification ratios that highlight our TR generality.
Benchmark Tokens Avg
64 96 128 192
TextVQA 49.6 52.9 54.2 55.7 53.1
[…]
POPE 57.3 71.7 77.3 82.1 72.1
[…]
5.3 Efficiency Analysis
Table 4: Efficiency analysis of LLaVA with SparseVLM. The detailed metric includes storage (cache memory), latency (CUDA time), and computation (FLOPs). $\Delta$ denotes the reduction ratio.
Method Token Acc. Storage $\Delta$ CUDA $\Delta$ FLOPs $\Delta$
Memory (MB) Time (ms) $\downarrow$ (T) $\downarrow$
Baseline 576 100% - - 419.9 - 9.6 -
FastV 192 88% 100.8 66.7% 290.8 30.7% 2.3 76.0%
SparseVLM 64 87% 33.6 88.9% 193.5 53.9% 1.5 84.4%
SparseVLM affords significant efficiency and storage gains for the inference process. We conduct a comparative analysis of storage memory, CUDA time, and FLOPs on LLaVA-7B, and compare our method with the baseline method and FastV (Chen et al., 2024b). As displayed in Table 4, we conduct an inference efficiency analysis on a single NVIDIA A100-80GB with identical lengths of text prompts and single-image inputs. Compared to the baseline model, SparseVLM achieves a significant reduction of $53.9\%$ in CUDA time and $84.4\%$ in FLOPs while keeping $88\%$ accuracy. Despite SparseVLM has an additional overhead to calculate text raters and cluster-pruned vision tokens, it leads to fewer than FastV tokens with comparable accuracy. Additionally, SparseVLM demonstrates lower metrics in terms of CUDA latency time and FLOPs by $23.2\%$ and $8.4\%$, respectively.
5.4 Visualization
As shown in Figure 6, we visualize SparseVLM on various VQA questions. From left to right, we visualize the results after we apply token pruning to different layers. On the very right, the dialogue box contains the prompt with the highlighted in red the most relevant text, and below is the generated answer in green produced by the remaining pruned image tokens. As the number of layers increases, more tokens are pruned and the Region of Interest (ROI) is gradually refined. The model systematically reduces less relevant image information while retaining key tokens that are closely tied to the question. The visualization reveals that SparseVLM, although discarding some overall image details, effectively retains essential visual tokens. These preserved tokens encapsulate the features necessary for answering the question, focusing on more relevant visual regions through their interaction with the question. More visualization cases can be found in the Appendix A.6.
6 Conclusion
This paper introduced an efficient text-aware training-free token optimization mechanism called SparseVLM which significantly enhanced the efficiency of various VLMs in image and video understanding tasks. Unlike existing methods, SparseVLM was able to optimize VLMs without introducing extra parameters and fine-tuning costs. We achieved a more compact visual representation by progressively pruning the less relevant vision tokens. In addition, we employed the matrix rank to adaptively determine sparsification ratios and recycled the pruned tokens via reconstruction to reduce the information loss. Experiments demonstrated that SparseVLM increased the efficiency of various VLMs. Particularly, LLaVA equipped with SparseVLM achieved a reduction of 53.9% latency with a compression ratio of 88.9% while maintaining 87% accuracy. Moreover, our method exceeded FastV accuracy by 34.4% in video understanding tasks. Our SparseVLM can provide practical benefits for deploying off-the-shelf VLMs on edge devices and in the cloud.
Appendix A Appendix
A.1 The Redundancy in Vision Tokens
In text-unrelated tasks, such as classification or segmentation, it is common to use a downsampling strategy which reduces redundancy in visual modality and makes the model more efficient to train (Zhang et al., 2024a). In Figure 7, which starts by comparing the original image with a downsampled version. The downsampled image reduces the number of tokens from 1166 to 576, achieving a 50% increase in efficiency. However, this process results in a 15% loss of information, as indicated by the decrease in entropy from 7.44 to 6.13. This trade-off is deemed acceptable for tasks unrelated to text such as classification or segmentation. For text-related tasks, such as visual question answering (VQA), there are two different modalities, text and vision. In this figure, the prompt is ”What is written on the top of the yellow sticker on the fridge?” The output generated is ”Warning”. Pay attention to the highlighted part in both text and image, the text with the highest information density is highlighted with color, accounting for 88% of the total text; the region of interest (related to the prompt) part in the image only rates 38% in the whole image, which demonstrates that the information in images is typically more sparse than in natural language. Therefore, we proposed the SparseVLM to prune redundancy in visual tokens progressively. With our method, visual redundancy is reduced while maintaining the essential information required for accurate task performance, effectively improving the model’s efficiency and effectiveness across different vision tasks.
Figure 7: Comparison of visual redundancy in different vision tasks.
A.2 Computing Budget Detailed Estimation
@@
Estimation of Visual Token Significance. In this stage, only the equation 4 averaging process requires computation. Each vision token undergoes $L_{t}-1$ additions and one division. With $L_{v}$ vision tokens in total, the number of FLOPs for this stage is $(L_{t}-1+1)\times L_{v}=L_{t}\times L_{v}$.
@@
A.3 Efficiency Details
We present a comparative efficiency analysis of SparseVLM, the baseline, and FastV (Chen et al., 2024b) during the inference phase in Table 4. In this section, we provide additional details on the CUDA time measurement during the inference phase. Following VoCo-LLaMA Ye et al. (2024) setting, we primarily consider the following components that contribute to the reported CUDA time: image encoding time (if applicable), kv cache load time (if applicable), and transformers forward time. We exclude other computational times that are not dependent on the model itself and the caching strategy, such as model loading time, from the CUDA time measurement. Specifically, the attention operation is implemented by Sdpa Attention: https://pytorch.org/tutorials/intermediate/scaled_dot_product_attention_tutorial.
@@ A.3 Efficiency Details
ConBench. (Zhang et al., 2024b) The ConBench benchmark predominantly focuses on the consistency of the model’s answers across a wide variety of different tasks and question types. It presents three core capabilities in a hierarchical manner, namely observation ability (sensation), complex reasoning (reasoning), and professional knowledge (knowledge). This hierarchical design aims to gradually challenge the performance of models on different tasks and provides fine-grained evaluation indicators, so as to evaluate the performance and consistency of the model.
@@
MSVD-QA. (Xu et al., 2017) The MSVD-QA benchmark is based on the existing Microsoft Research Video Description (MSVD) dataset and contains 1970 video clips and approximately 50.5K QA pairs. The questions and answers are diverse in nature, covering a wide range of topics and aspects related to the video content. Due to its relatively large data size and the diversity of questions, it is widely used for video question answering tasks and video caption tasks. The tasks formed in it are open-ended questions, consisting of five types of questions: what, who, how, when and where.
@@
A.5 Implementation Details.
All of our experiments are conducted on a single Nvidia A100-80G GPU. The implementation was carried out in Python 3.10, utilizing PyTorch 2.1.2, CUDA 11.8, and transformers 4.31.0. The inference follows the evaluation settings established by LLaVA(Liu et al., 2024).
A.6 Visualization
Figure 8 contains more visualization examples of SparseVLM on various VQA prompts.

### v3

@@ A.2 Compatibility with FlashAttention
    Here, each block $B$ is processed independently, and the results are combined using incremental normalization.
A.3 Computing Budget Detailed Estimation
@@
Estimation of Visual Token Significance. In this stage, only the equation 3 averaging process requires computation. Each visual token undergoes $L_{t}-1$ additions and one division. With $L_{v}$ visual tokens in total, the number of FLOPs for this stage is $(L_{t}-1+1)\times L_{v}=L_{t}\times L_{v}$.
@@ A.4 Dataset
MMVet. (Yu et al., 2023) The MMVet benchmark is designed based on the insight that the intriguing ability to solve complicated tasks is often achieved by a generalist model being able to integrate different core vision-language capabilities. MM-Vet defines 6 core VL capabilities and examines the 16 integrations of interest derived from the capability combination.
@@
MSVD-QA. (Xu et al., 2017) The MSVD-QA benchmark is based on the existing Microsoft Research Video Description (MSVD) dataset and contains 1970 video clips and approximately 50.5K QA pairs. The questions and answers are diverse in nature, covering a wide range of topics and aspects related to the video content. Due to its relatively large data size and the diversity of questions, it is widely used for video question answering tasks and video caption tasks. The tasks formed in it are open-ended questions, consisting of five types of questions: what, who, how, when, and where.
@@
A.5 Implementation Details
All of our experiments are conducted on a single NVIDIA A100-80G GPU. The implementation is carried out in Python 3.10, utilizing PyTorch 2.1.2, CUDA 11.8, and transformers 4.31.0. The inference follows the evaluation settings established by LLaVA (Liu et al., 2024). For LLaVA-1.5-7/13B, Mini-Gemini (MGM), and Qwen-VL, we follow the same inference setting as the original paper as it is publicly available111github.com/haotian-liu/LLaVA 222github.com/dvlab-research/MGM  [^3] https://github.com/QwenLM/Qwen-VL. For video understanding tasks, we adopt the same inference setup as the original Video-LLaVA code base444github.com/PKU-YuanGroup/Video-LLaVA., as it is publicly available.
A.6 Efficiency Details
We present a comparative efficiency analysis of SparseVLM, the baseline, and FastV (Chen et al., 2024) during the inference phase in Table 1. In this section, we provide additional details on the CUDA time during the inference phase. Following VoCo-LLaMA (Ye et al., 2024), we primarily consider the following components that contribute to the reported CUDA time: image encoding time (if applicable), KV cache load time (if applicable), and transformers forward time. We exclude other computational times that are not dependent on the model itself and the caching strategy, such as model loading time, from the CUDA time measurement. Specifically, the attention operation is implemented by FlashAttention (Dao et al., 2022).
A.7 More Sparsification Visualization
Figure 8 showcases a diverse array of visualization examples that demonstrate the application of SparseVLM across a spectrum of Visual Question Answering (VQA) prompts. These visualizations offer a deeper insight into how SparseVLM processes and responds to different types of queries posed in a visual context.
@@ A.7 More Sparsification Visualization
A.8 More Detailed Efficiency Analysis
To better validate the efficiency of our method, we provide the latency-vs.-accuracy and FLOPs-vs.-Accuracy trade-offs for SparseVLM applied to LLaVA and MGM across three benchmarks: POPE, TextVQA, and MME, which are shown in Figure 9 and Figure 10. Besides, we also analyze Video-LLaVA matched with SparseVLM in Figure 11 on TGIF and MSVD.
Figure 9: Trade-offs for SparseVLM on LLaVA: (a) Latency vs. Accuracy, and (b) FLOPs vs. Accuracy. Both show comparisons of random sparse, SparseVLM, and baseline.
Figure 10: Trade-offs for SparseVLM on MGM: (a) Latency vs. Accuracy, and (b) FLOPs vs. Accuracy. Both show comparisons of random sparse, SparseVLM, and baseline.
Figure 11: Trade-offs for SparseVLM on Video-LLaVA: (a) Latency vs. Accuracy, and (b) Token budget vs. Accuracy. Both show comparisons of SparseVLM, FastV, and baseline.